–6 + 7(x – 1) = 2x + 5(x – 6) – 4
Answer by checkley77(12844)

–6 + 7(x – 1) = 2x + 5(x – 6) – 4
-6+7x-7=2x+5x-30-4
7x-2x-5x=-30-4+6+7
0x=21 no solution for x.